# Multiple quantum exceptional, diabolical, and hybrid points in multimode bosonic systems: II. Nonconventional $\mathcal{PT}$-symmetric dynamics and unidirectional coupling

### v1

@@ Abstract
We analyze the existence and degeneracies of quantum exceptional, diabolical, and hybrid points of simple bosonic systems, composed of up to six modes with damping and/or amplification and exhibiting nonconventional dynamics. They involve the configurations in which the dynamics typical for PT-symmetric systems is observed only in a subspace of the whole Liouville space of the system states (nonconventional PT-symmetric dynamics) as well as those containing unidirectional coupling. The system dynamics described by quadratic non-Hermitian Hamiltonians is governed by the Heisenberg-Langevin equations. Conditions for the observation of inherited quantum hybrid points with up to sixth-order exceptional and second-order diabolical degeneracies are revealed, though relevant only for short-time dynamics. This raises the question of whether higher-order inherited singularities exist in bosonic systems that exhibit physically meaningful behavior at arbitrary times. On the other hand, for short times, unidirectional coupling of various types enables the concatenation of simple bosonic systems with second- and third-order exceptional degeneracies on demand. This approach allows for the creation of arbitrarily high exceptional degeneracies observed in systems with diverse structures. Methods for numerical identifying the quantum exceptional and hybrid points, and determining their degeneracies are discussed. Rich dynamics of higher-order field-operator moments is analyzed from the point of view of the presence of exceptional and diabolical points with their degeneracies in general.
I Introduction
@@
Non-Hermitian bosonic $\mathcal{PT}$-symmetric systems are endowed with interesting physical properties that occur at and around quantum exceptional and hybrid points (QEPs and QHPs). Some studies [1, 2, 3, 4, 5] suggest that these points can enable improvements in measurement precision beyond the classical limit. They also allow to enhance effective system nonlinearities that result in the generation of highly nonclassical and entangled states [6, 7]. The strength of these effects depends in many cases on the order of the degeneracy of exceptional points (EPs): The higher-order the degeneracy, the more enhanced the processes become. This lead us in part I of the paper [8] to the analysis of QEPs and QHPs of parity-time (PT) -symmetric bosonic systems with up to five modes considering different configurations. However, this analysis revealed only the systems with QHPs with the second- and third-order exceptional geneneracies (EDs) and second-order diabolical degeneracies (DDs), despite the fact that the bosonic systems with five modes were considered with the promise of observation of QHPs with the fifth-order ED.
@@
For this reason, we extend our previous analysis [8] in two directions to allow for the observation of higher-order EDs for inherited QEPs and QHPs. First, we weaken our requirements for the observation of $\mathcal{PT}$-symmetric dynamics by considering only subspace(s) of the whole Liouville space of the statistical operators. We note here that, owing to the linearity of quantum mechanics, we can equivalently describe the system dynamics [9, 10] in the Liouville space of the statistical operators and the complete space spanned by the operators of measurable quantities. We also note that the linearity gives the one-to-one correspondence between the subspaces of the above mentioned spaces. When such $\mathcal{PT}$-symmetric-like behavior is restricted to only a subspace we refer to nonconventional $\mathcal{PT}$-symmetric system behavior.
@@
Second, we admit in our analysis more general non-Hermitian $\mathcal{PT}$symmetric Hamiltonians. We recall here that, in Ref. [8], the non-Hermiticity of the investigated Hamiltonians originated only in the presence of damping and amplification terms, whose non-Hermiticity was ‘remedied’ by the presence of the Langevin fluctuating operator forces [11, 10]. This guarantees the system evolution preserving the bosonic canonical commutation relations. Here, we consider also the Hamiltonians that describe bosonic systems with unidirectional coupling between the modes. The reason is that unidirectional coupling allows to concatenate two bosonic subsystems such that they keep their original eigenvalues. Moreover the original subspaces belonging to the same eigenvalues merge together which results in the increased EDs. This property gives rise to the method suggested and elaborated in Refs. [12, 13] that provides QEPs with higher-order EDs. However, we note that, apart from a complex experimental realization, unidirectional coupling is highly non-Hermitian and violates reciprocity of physical processes. Nevertheless, up to our best knowledge, this is the only straightforward method for reaching QEPs and QHPs with high-order EDs for open bosonic systems described correctly according to quantum mechanics. Higher-order QEPs using unidirectional coupling were already realized in [14].
@@
We note that when Hamiltonians are directly considered, their higher-order EPs can relatively easily be observed. For example, higher-order EPs were predicted in optomechanical [15] and cavity magnonic systems [16], those described by the Bose-Hubbard model [17], or photonic structures [18, 19]. Higher-order EDs were also studied in Refs. [20, 21]. They play significant role in amplification [22] and sensing [23] as well as speeding up entanglement generation [24].
@@
Increasing complexity of the bosonic systems poses the question about identification of inherited QEPs and QHPs and the determination of their degeneracies. Whereas simple bosonic systems allow for analytical derivation of the eigenvalues and eigenvectors of their dynamical matrices, more complex bosonic systems admit only numerical treatment. In this case, we may numerically decompose a given dynamical matrix into its Jordan form that directly reveals QEPs and QHPs with their degeneracies. Alternatively, we may add a little $\epsilon$ perturbation to any element of the dynamical matrix that can remove both EDs and DDs and even allow for distinguishing ED and DD.
@@
Some of the effects related to the presence of QEPs and QHPs with higher-order EDs and DDs are observed also in the behavior of higher-order field-operator moments (FOMs) [11, 9, 25]. We note that we refer to genuine QEPs and QHPs in the case of higher-order FOMs. This originates from the fact that the dynamics of $n$th-order FOMs is built, in certain sense, as a ‘multiplied dynamics’ of first-order FOMs with its inherited QEPs and QHPs. This ‘multiplied dynamics’ then naturally contains the genuine QEPs and QHPs with higher-order EDs and DDs, as it was discussed in [25]. In general, genuine QEPs with ED orders up to $n$th power of ED orders of inherited QEPs of the first-order FOMs are expected in the dynamics of $n$th-order FOMs. However, the structure of genuine QEPs and QHPs in the dynamics of higher-order FOMs intimately depends on that of the inherited QEPs and QHPs found for the first-order FOMs. For this reason, we have explicitly revealed the structure of genuine QEPs and QHPs belonging to the second-order FOMs for the bosonic systems in tables of Ref. [8]. They explicitly elucidate the relation between EDs and DDs of these genuine QEPs and QHPs and degeneracies of inherited QEPs and QHPs. We note that induced QEPs and QHPs, which were introduced in Ref. [25], have their origin in the existence of identical or similar (related by the field commutation relations) FOMs in the formal construction of higher-order FOMs spaces and they further increase the multiplicity of spectral degeneracies. Nevertheless, they do not enrich the system dynamics. There exist some general properties of EDs and DDs of genuine QEPs and QHPs of bosonic systems independent of their configuration that we address here to complete the analysis of specific bosonic systems.
@@
The paper is organized as follows. Section II is devoted to bosonic systems exhibiting nonconventional $\mathcal{PT}$-symmetric dynamics, An an example, four-mode systems are analyzed. Section III contains the analysis of bosonic systems with unidirectional coupling in linear configurations involving in turn from two to six modes. The dynamics of the two-mode bosonic system with unidirectional coupling and its applicability are analyzed in Sec. IV. Numerical methods for the identification of QEPs and QHPs and their degeneracies are discussed in Sec. V. A general analysis of genuine and induced QHPs in the dynamics of arbitrary-order FOMs is given in Sec. VI. Section VII brings conclusions. In Appendix A, the properties of the Langevin operator forces are discussed considering the model with unidirectional coupling. Statistical properties of the two-mode bosonic system with unidirectional coupling are described in Appendix B.
@@
II Bosonic systems with bidirectional coupling and nonconventional $\mathcal{PT}$-symmetric dynamics
@@
Figure 1: Schematic diagrams of the four-mode bosonic systems in (a) circular and (b) tetrahedral configurations that exhibit QHPs in their nonconventional $\mathcal{PT}$-symmetric dynamics. Strengths $\epsilon$ and $\kappa$ characterize, respectively, the linear and nonlinear coupling between the modes, while $\gamma$, with subscripts indicating the mode number(s), are the damping or amplification rates, and the annihilation operators $\hat{a}$ identify the mode number via their subscripts.
II.1 Circular configuration
@@
[…]
where $\hat{a}_{j}$ ($\hat{a}_{j}^{\dagger}$) for $j=1,\ldots,4$ denotes the annihilation (creation) operator of the $j$th mode, $\epsilon$ ($\kappa$) is the linear (nonlinear) coupling strength between the modes [26]. Symbol H.c. replaces the Hermitian-conjugated terms. Damping or amplification of mode $j$ is described by the damping (amplification) rate $\gamma_{j}$ and the corresponding Langevin stochastic operator forces, $\hat{L}_{j}$ and $\hat{L}_{j}^{\dagger}$ that occur in the dynamical Heisenberg-Langevin equations written below in Eq. (2). The Langevin stochastic operator forces are assumed to have the Markovian and Gaussian properties specific to the damping and amplification processes [25]. Their presence in the Heisenberg-Langevin equations guarantees the fulfillment of the field-operator commutation relations. Moreover the properties of the Langevin stochastic operator forces are related to the damping (amplification) rates via the fluctuation-dissipation theorems [10, 11].
@@
where the vectors $\hat{\bm{a}}$ of field operators and $\hat{\bm{L}}$ of the Langevin operator forces are given as $\hat{\bm{a}}=\left[\hat{\bm{a}}_{\bm{1}},\hat{\bm{a}}_{\bm{2}},\hat{\bm{a}}_{\bm{3}},\hat{\bm{a}}_{\bm{4}}\right]^{T}\equiv\left[\hat{a}_{1},\hat{a}_{1}^{\dagger},\hat{a}_{2},\hat{a}_{2}^{\dagger},\hat{a}_{3},\hat{a}_{3}^{\dagger},\hat{a}_{4},\hat{a}_{4}^{\dagger}\right]^{T}$ and $\hat{\bm{L}}=\left[\hat{L}_{1},\hat{L}_{1}^{\dagger},\hat{L}_{2},\hat{L}_{2}^{\dagger},\hat{L}_{3},\hat{L}_{3}^{\dagger},\hat{L}_{4},\hat{L}_{4}^{\dagger}\right]^{T}$. The dynamical matrix $\bm{M^{(4)}_{\rm c}}$ introduced in Eq. (2) is derived in the form
@@
Applying the conditions
@@
in the dynamical matrix $\bm{M^{(4)}_{\rm c}}$ in Eq. (7), we reveal its eigenvalues $\lambda_{j}^{M^{(4)}_{\rm c}}$:
@@
If $\beta=0$ then $\lambda_{3}^{M^{(4)}_{\rm c}}=\lambda_{4}^{M^{(4)}_{\rm c}}$ and also $\bm{y_{3}^{M^{(4)}_{\rm c}}}=\bm{y_{4}^{M^{(4)}_{\rm c}}}$. We note that the eigenvalues $\lambda_{3}^{M^{(4)}_{\rm c}}$ and $\lambda_{4}^{M^{(4)}_{\rm c}}$ share also their imaginary parts, which is important for the observation of $\mathcal{PT}$-symmetric-like dynamics in a suitable interaction frame [27]. Provided that the system initial conditions are chosen such that only the eigenvalues $\lambda_{3}^{M^{(4)}_{\rm c}}$ and $\lambda_{4}^{M^{(4)}_{\rm c}}$ determine its dynamics, we observe a second-order QEP. This QEP changes into a QHP with second-order ED and DD when the $8\times 8$ matrix $\bm{M^{(4)}_{\rm c}}$ is analyzed. The condition $\beta=0$ transforms into the formula
@@
(a) Refer to caption
(b) Refer to caption
Figure 2: Real parts $\lambda^{\rm r}$ of the eigenvalues (a) $\lambda_{3,4}^{M^{(4)}_{\rm c}}$ of the matrix $\bm{M^{(4)}_{\rm c}}$, given in Eq. (7), for the four-mode bosonic system in the circular configuration with different damping and/or amplification rates of neighbor modes and (b) $\lambda_{3,4}^{M^{(4)}_{\rm t}}$ of the matrix $\bm{M^{(4)}_{\rm t}}$, given in Eq. (25), for $\xi=\pm\zeta$ for the four-mode bosonic system in the tetrahedral configuration with the same damping and/or amplification rates of neighbor modes are drawn in the parameter space $(\kappa/\epsilon,\gamma_{-}/\epsilon)$. Dashed red curves indicate the positions of the QHPs and are given by Eq. (16).

### v2

@@ Abstract
We analyze the existence and degeneracies of quantum exceptional, diabolical, and hybrid points in simple bosonic systems - comprising up to six modes with damping and/or amplification - under two complementary scenarios to those of Ref. [1]: (i) nonconventional PT-symmetric dynamics confined to a subspace of the full Liouville space, and (ii) systems featuring unidirectional coupling. The system dynamics described by quadratic non-Hermitian Hamiltonians is governed by the Heisenberg-Langevin equations. Conditions for the observation of inherited quantum hybrid points with up to sixth-order exceptional and second-order diabolical degeneracies are revealed, though relevant only for short-time dynamics. This raises the question of whether higher-order inherited singularities exist in bosonic systems under general conditions. Nevertheless, for short times, unidirectional coupling of various types enables the concatenation of simple bosonic systems with second- and third-order exceptional degeneracies such that arbitrarily high exceptional degeneracies are reached. Methods for numerical identifying the quantum exceptional and hybrid points together with their degeneracies are addressed. Following Ref. [1] rich dynamics of second-order field-operator moments is analyzed from the point of view of the presence of exceptional and diabolical points and their degeneracies.
1 Introduction
Non-Hermitian bosonic parity-time ($\mathcal{PT}$) -symmetric systems exhibit a range of remarkable phenomena at and near quantum exceptional and hybrid points (QEPs and QHPs, respectively). Several studies [2, 3, 4, 5, 6, 7] have shown that these singularities can enhance measurement precision beyond the classical limit and amplify effective system nonlinearities, leading to the generation of highly nonclassical and entangled states with tailored properties [8, 9, 10, 11]. We note that several studies (e.g., [12, 13, 14, 15, 16, 17]) have demonstrated that in the linear regimes of quantum systems the measurement sensitivity boost may not occur. At present, investigations of QEPs in nonlinear quantum systems from the point of view of the measurement precision proceed to demonstrate the advantage of $\mathcal{PT}$-symmetric dynamics in connection with the system nonlinearity that allows on its own to beat the classical limit. However, the intrinsic noise accompanying damping and amplification in $\mathcal{PT}$-symmetric quantum systems must also be taken into account, as it typically degrades quantumness over extended timescales [18, 19]. Despite this, highly squeezed and sub-Poissonian states remain achievable. Under various configurations of passive and active $\mathcal{PT}$-symmetric bosonic systems, one can also generate entangled states exhibiting asymmetric steering and Bell nonlocality [20]. Moreover, these systems — and their field-operator moments (FOMs) [21] — can simulate complex many-body dynamics, including boundary effects [21, 22]. On the applied side, optical switching via encirclement of a QHP has been demonstrated [23], and unidirectional light propagation — along with invisibility cloaking — has been realized in such platforms [24, 25, 26, 27]. For a comprehensive overview of the rich physics around the spectral singularities of $\mathcal{PT}$-symmetric systems, see the reviews in [28, 29].
@@ 1 Introduction
The strength of these effects depends in many cases on the order of the degeneracy of exceptional points (EPs): The higher-order the degeneracy, the more enhanced the processes become. This led us in Part I of Ref. [1] to the analysis of QEPs and QHPs of $\mathcal{PT}$-symmetric bosonic systems with up to five modes considering different configurations under very general conditions. However, this analysis revealed only the systems with QHPs with the second- and third-order exceptional degeneracies (EDs) and second-order diabolical degeneracies (DDs), despite the fact that the bosonic systems with five modes were considered with the promise of observation of QHPs with the fifth-order ED.
@@
For this reason, we consider more general bosonic system compared to those analyzed in [1] to seek for the observation of higher-order EDs for inherited QEPs and QHPs. First, we weaken our requirements for the observation of $\mathcal{PT}$-symmetric dynamics by considering only subspace(s) of the whole Liouville space of the statistical operators. We note here that, owing to the linearity of quantum mechanics, we can equivalently describe the system dynamics [30, 31] in the Liouville space of the statistical operators and the complete space spanned by the operators of measurable quantities. We also note that the linearity gives the one-to-one correspondence between the subspaces of the above mentioned spaces. When such $\mathcal{PT}$-symmetric-like behavior is restricted to only a subspace we refer to nonconventional $\mathcal{PT}$-symmetric system behavior.
@@
Second, we admit in our analysis more general non-Hermitian $\mathcal{PT}$ -symmetric Hamiltonians. We recall here that, in Ref. [1], the non-Hermiticity of the investigated Hamiltonians originated only in the presence of damping and amplification terms, whose non-Hermiticity was ‘remedied’ by the presence of the Langevin fluctuating operator forces [32, 31]. This guarantees the system evolution preserving the bosonic canonical commutation relations. Here, we consider also the Hamiltonians that describe bosonic systems with unidirectional coupling between the modes. The reason is that unidirectional coupling allows to concatenate two bosonic subsystems such that they keep their original eigenvalues. Moreover the original subspaces belonging to the same eigenvalues merge together which results in the increased EDs. This property gives rise to the method suggested and elaborated in Refs. [33, 34] that provides QEPs with higher-order EDs. Realization of unidirectional propagation is based upon ring resonators  [33, 34]. However, we note that, apart from a complex experimental realization, unidirectional coupling is highly non-Hermitian and violates reciprocity of physical processes. Nevertheless, up to our best knowledge, this is the only method for reaching QEPs and QHPs with high-order EDs for open bosonic systems under general conditions. However, as we show here, the quantum-mechanical consistency of the models with unidirectional coupling is guaranteed only for short times. Higher-order QEPs using unidirectional coupling were already realized in [35].
@@
We note that if quantum field-amplitude fluctuations are completely neglected and pure states are considered, higher-order EPs in the Hamiltonian spectra can be observed relatively easily. For example, higher-order EPs were predicted in optomechanical [36] and cavity magnonic systems [37], those described by the Bose-Hubbard model [38], or photonic structures [39, 40]. Higher-order EDs were also studied in Refs. [41, 42]. They play significant role in amplification [43] and sensing [5] as well as speeding up entanglement generation [44]. However, the dynamics of open quantum systems brings new features and qualitatively modifies the systems’ behavior [45, 46].
@@
Increasing complexity of the bosonic systems also poses the question about identification of inherited QEPs and QHPs and the determination of their degeneracies. Whereas simple bosonic systems allow for analytical derivation of the eigenvalues and eigenvectors of their dynamical matrices, more complex bosonic systems admit only numerical treatment. In this case, we may numerically decompose a given dynamical matrix into its Jordan form that directly reveals QEPs and QHPs with their degeneracies. Alternatively, we may add a little $\epsilon$ perturbation to any element of the dynamical matrix that can remove both EDs and DDs and even allow for distinguishing ED and DD.
@@
Some of the effects related to the presence of QEPs (quantum exceptional points) and QHPs (quantum hybrid points) with higher-order EDs (exceptional degeneracies) and DDs (diabolical degeneracies) are observed also in the behavior of higher-order FOMs (field-operator moments) [32, 30, 47]. We note that we refer to genuine QEPs and QHPs in the case of higher-order FOMs. This originates from the fact that the dynamics of $n$th-order FOMs is built, in certain sense, as a ‘multiplied dynamics’ of first-order FOMs with their inherited QEPs and QHPs. This ‘multiplied dynamics’ then naturally contains the genuine QEPs and QHPs with higher-order EDs and DDs, as it was discussed in [47]. In general, genuine QEPs with ED orders up to $n$th power of ED orders of inherited QEPs of the first-order FOMs are expected in the dynamics of $n$th-order FOMs (for examples, see Ref. [47]). The structure of genuine QEPs and QHPs in the dynamics of higher-order FOMs intimately depends on that of the inherited QEPs and QHPs found for the first-order FOMs. For this reason, we have explicitly revealed the structure of genuine QEPs and QHPs belonging to the second-order FOMs for the bosonic systems in the tables of Appendix B of Ref. [1]. They explicitly elucidate the relation between EDs and DDs of these genuine QEPs and QHPs and degeneracies of inherited QEPs and QHPs. We note that the induced QEPs and QHPs, which were introduced in Ref. [47], have their origin in the existence of identical or similar (related by the field commutation relations) FOMs in the formal construction of higher-order FOMs spaces and they further increase the multiplicity of spectral degeneracies. Nevertheless, they do not enrich the system dynamics. There exist some general properties of EDs and DDs of genuine QEPs and QHPs of bosonic systems independent of their configuration that we address here to complete the analysis of specific bosonic systems.
@@
The paper is organized as follows. Section 2 is devoted to bosonic systems exhibiting nonconventional $\mathcal{PT}$-symmetric dynamics, as an example, four-mode systems are analyzed. Section 3 contains the analysis of bosonic systems with unidirectional coupling in linear configurations involving in turn from two to six modes. The dynamics of the two-mode bosonic system with unidirectional coupling and its applicability are analyzed in Sec. 4. A general analysis of genuine and induced QHPs in the dynamics of arbitrary-order FOMs is given in Sec. 5. Section 6 brings conclusions. Tables summarizing QEPs and QHPs—along with their eigenvalue EDs and DDs—for first- and second-order FOM dynamics are given in Appendix A. Appendix B presents the numerical methods used to identify these singularities and their degeneracies. In Appendix C, we analyze the properties of the Langevin operator forces in the unidirectional-coupling model, while Appendix D details the statistical characteristics of the two-mode bosonic system under unidirectional coupling.
@@
2 Bosonic systems with bidirectional coupling and nonconventional $\mathcal{PT}$-symmetric dynamics
@@
Figure 1: Schematic diagrams of the four-mode bosonic systems in (a) circular and (b) tetrahedral configurations that exhibit quantum exceptional points (QEPs) and quantum hybrid points (QHPs) with various exceptional degeneracies (EDs) and diabolical degeneracies ($D$s) observed in their nonconventional $\mathcal{PT}$-symmetric dynamics of field-operator moments (FOMs) of different orders. Strengths $\epsilon$ and $\kappa$ characterize, respectively, the linear and nonlinear coupling between the modes, while $\gamma$, with subscripts indicating the mode number(s), are the damping or amplification rates, and the annihilation operators $\hat{a}$ identify the mode number via their subscripts.
2.1 Circular configuration
@@
$\left.+\hbar\kappa\hat{a}_{2}\hat{a}_{3}+\hbar\kappa\hat{a}_{3}\hat{a}_{4}+\hbar\kappa\hat{a}_{4}\hat{a}_{1}\right]+\textrm{H.c.}$
where $\hat{a}_{j}$ ($\hat{a}_{j}^{\dagger}$) for $j=1,\ldots,4$ denotes the annihilation (creation) operator of the $j$th mode, $\epsilon$ ($\kappa$) is the linear (nonlinear) coupling strength between the modes [48]. Symbol H.c. replaces the Hermitian-conjugated terms. Damping or amplification of mode $j$ is described by the damping (amplification) rate $\gamma_{j}$ and the corresponding Langevin stochastic operator forces, $\hat{L}_{j}$ and $\hat{L}_{j}^{\dagger}$ that occur in the dynamical Heisenberg-Langevin equations written below in Eq. (2). The Langevin stochastic operator forces are assumed to have the Markovian and Gaussian properties specific to the damping and amplification processes [47]. In Eq. (98), we present an example of first- and second-order correlation functions of the Langevin forces associated with damping (in mode 1) and amplification (in mode 2). Their presence in the Heisenberg-Langevin equations guarantees the fulfillment of the field-operator commutation relations. Moreover the properties of the Langevin stochastic operator forces are related to the damping (amplification) rates via the fluctuation-dissipation theorems [31, 32].
@@
where the vectors $\hat{\bm{a}}$ of field operators and $\hat{\bm{L}}$ of the Langevin operator forces are given as $\hat{\bm{a}}=\left[\hat{\bm{a}}_{\bm{1}},\hat{\bm{a}}_{\bm{2}},\hat{\bm{a}}_{\bm{3}},\hat{\bm{a}}_{\bm{4}}\right]^{T}\equiv\left[\hat{a}_{1},\hat{a}_{1}^{\dagger},\hat{a}_{2},\hat{a}_{2}^{\dagger},\hat{a}_{3},\hat{a}_{3}^{\dagger},\hat{a}_{4},\hat{a}_{4}^{\dagger}\right]^{T}$ and $\hat{\bm{L}}=\left[\hat{L}_{1},\hat{L}_{1}^{\dagger},\hat{L}_{2},\hat{L}_{2}^{\dagger},\hat{L}_{3},\hat{L}_{3}^{\dagger},\hat{L}_{4},\hat{L}_{4}^{\dagger}\right]^{T}$. The dynamical $8\times 8$ matrix $\bm{M^{(4)}_{\rm c}}$ introduced in Eq. (2) is derived in the form
@@
The submatrices $\bm{\tilde{\gamma}_{j}}$, $j=1,\dots,4$, and $\bm{\xi}$ in Eq. (12) can simultaneously be diagonalized. As the submatrices $\bm{\tilde{\gamma}_{j}}$ are linearly proportional to the identity matrix, the appropriate transformation does not change their form. On the other hand, it yields two eigenvalues $\lambda^{\xi}_{1,2}$ of the matrix $\bm{\xi}$ which we identify by a new variable $\xi$. In this diagonalized form of the above submatrices, the original $8\times 8$ dynamical matrix $\bm{M^{(4)}_{\rm c}}$ attains the form
$\bm{M^{(4)}_{\rm c}}=\left[\begin{array}[]{cccccccc}{\color[rgb]{0,0,1}-i{\gamma}_{1}}&0&{\color[rgb]{0,0,1}\lambda^{\xi}_{1}}&0&{\color[rgb]{0,0,1}0}&0&{\color[rgb]{0,0,1}\lambda^{\xi}_{1}}&0\\ 0&{\color[rgb]{0,1,0}-i{\gamma}_{1}}&0&{\color[rgb]{0,1,0}\lambda^{\xi}_{2}}&0&{\color[rgb]{0,1,0}0}&0&{\color[rgb]{0,1,0}\lambda^{\xi}_{2}}\\ {\color[rgb]{0,0,1}\lambda^{\xi}_{1}}&0&{\color[rgb]{0,0,1}-i{\gamma}_{2}}&0&{\color[rgb]{0,0,1}\lambda^{\xi}_{1}}&0&{\color[rgb]{0,0,1}0}&0\\ 0&{\color[rgb]{0,1,0}\lambda^{\xi}_{2}}&0&{\color[rgb]{0,1,0}-i{\gamma}_{2}}&0&{\color[rgb]{0,1,0}\lambda^{\xi}_{2}}&0&{\color[rgb]{0,1,0}0}\\ {\color[rgb]{0,0,1}0}&0&{\color[rgb]{0,0,1}\lambda^{\xi}_{1}}&0&{\color[rgb]{0,0,1}-i{\gamma}_{3}}&0&{\color[rgb]{0,0,1}\lambda^{\xi}_{1}}&0\\ 0&{\color[rgb]{0,1,0}0}&0&{\color[rgb]{0,1,0}\lambda^{\xi}_{2}}&0&{\color[rgb]{0,1,0}-i{\gamma}_{3}}&0&{\color[rgb]{0,1,0}\lambda^{\xi}_{2}}\\ {\color[rgb]{0,0,1}\lambda^{\xi}_{1}}&0&{\color[rgb]{0,0,1}0}&0&{\color[rgb]{0,0,1}\lambda^{\xi}_{1}}&0&{\color[rgb]{0,0,1}-i{\gamma}_{4}}&0\\ 0&{\color[rgb]{0,1,0}\lambda^{\xi}_{2}}&0&{\color[rgb]{0,1,0}0}&0&{\color[rgb]{0,1,0}\lambda^{\xi}_{2}}&0&{\color[rgb]{0,1,0}-i{\gamma}_{4}}\end{array}\right].$ (21)
In this form, the matrix $\bm{M^{(4)}_{\rm c}}$ is decomposed into the direct sum of two $4\times 4$ matrices belonging to $\xi=\lambda^{\xi}_{1}$ and $\xi=\lambda^{\xi}_{2}$. This allows us to effectively consider the dynamical matrix $\bm{M^{(4)}_{\rm c}}$ in its $4\times 4$ block structure as if there are just numbers instead of $2\times 2$ submatrices. This effectively halves the dimension of the diagonalization procedure, allowing us to derive important analytical results.
To reveal QEPs and QHPs, inspired by $\mathcal{PT}$-symmetry and the results of Part I of Ref. [1], we apply the conditions
@@
in the dynamical $4\times 4$ matrix $\bm{M^{(4)}_{\rm c}}$ in Eq. (7). Then, we reveal its eigenvalues $\lambda_{j}^{M^{(4)}_{\rm c}}$:
@@
If $\beta=0$ then $\lambda_{3}^{M^{(4)}_{\rm c}}=\lambda_{4}^{M^{(4)}_{\rm c}}$ and also $\bm{y_{3}^{M^{(4)}_{\rm c}}}=\bm{y_{4}^{M^{(4)}_{\rm c}}}$. We note that the eigenvalues $\lambda_{3}^{M^{(4)}_{\rm c}}$ and $\lambda_{4}^{M^{(4)}_{\rm c}}$ share also their imaginary parts, which is important for the observation of $\mathcal{PT}$-symmetric-like dynamics in a suitable interaction frame [49]. Provided that the system initial conditions are chosen such that only the eigenvalues $\lambda_{3}^{M^{(4)}_{\rm c}}$ and $\lambda_{4}^{M^{(4)}_{\rm c}}$ determine its dynamics, we observe a second-order QEP. This QEP changes into a QHP with second-order ED and DD when the $8\times 8$ matrix $\bm{M^{(4)}_{\rm c}}$ is analyzed (see below). The condition $\beta=0$ assuming $\xi^{2}=\zeta^{2}=\epsilon^{2}-\kappa^{2}$ [see below Eq. (27)] transforms into the formula
@@
(a) Refer to caption (b) Refer to caption
Figure 2: Real parts $\lambda^{\rm r}$ of the eigenvalues (a) $\lambda_{3,4}^{M^{(4)}_{\rm c}}$ of the matrix $\bm{M^{(4)}_{\rm c}}$, given in Eq. (23), for the four-mode bosonic system in the circular configuration with different damping and/or amplification rates of neighbor modes and (b) $\lambda_{3,4}^{M^{(4)}_{\rm t}}$ of the matrix $\bm{M^{(4)}_{\rm t}}$, given in Eq. (47), for $\xi=\pm\zeta$ for the four-mode bosonic system in the tetrahedral configuration with the same damping and/or amplification rates of neighbor modes are drawn in the parameter space $(\kappa/\epsilon,\gamma_{-}/\epsilon)$. The dashed red curves indicate the values at the positions of the QHPs of the four-mode bosonic systems as given by Eq. (25).